a rectangular pool is 6m x 7.5m and is surrounded by a path 1m wide. what is the area of the path
Answer by dkppathak(439)

a rectangular pool is 6m x 7.5m and is surrounded by a path 1m wide. what is the area of the path
length of pool including path=6+2=8m
breath of pool including path =7.5+2=9.5 m
area of path= area of outer -are of inner =8x9.5-6x7.5= 76-45=31 M^2